Find the value of m for which the system of equations

x - 2y + z = 0,

-2x - y + 3z = 0

y + z = m 

has only trivial solution.


1
Expert's answer
2022-03-24T19:25:49-0400

Since trivial solution is (x,y,z)=0, we have m=0.

The system is

"\\left\\{ \\begin{array}{c}\tx-2y+z=0\\\\\t-2x-y+3z=0\\\\\ty+z=0\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tx-2y-y=0\\\\\t-2x-y-3y=0\\\\\tz=-y\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tx=3y\\\\\tx=-2y\\\\\tz=-y\\\\\\end{array} \\right. \\Rightarrow x=y=z=0"

Thus m=0 is the answer.